# Navigating Public Sentiment in the Circular Economy through Topic Modelling and Hyperparameter Optimisation

### v1

@@ 4.1.3 Twitter
In conclusion, BERTopic provided a deep extraction of CE topics through a large language model, CorEx provided a unique separation of themes (highest coherence score), and LDA captured the underlying concepts of CE on Twitter, all of which provide valuable insight into our understanding of the public sentiment on the CE. In the next three subsections, we will visualise and discuss the topics generated by the best models on each dataset in more depth.
@@
4.2 Official Sources: The Guardian
The remaining five topics each feature one or two words with significantly higher scores. For instance, in Topic 14, the score for ‘ev’ is notably higher than that of the other words. This is indicative of the strength of BERTopic as a large language model that utilises contextual analysis to effectively summarise and comprehend abbreviations like ‘ev’. Other related terms, such as brands including ‘Tesla’ and ‘Nissan’, aid the model in more swiftly grasping the significance of ‘ev’. The descending order of scores also illustrates the prioritisation and segmentation of these words within the topic.
We acknowledge that BERTopic performs admirably in interpreting text within The Guardian dataset, aggregating relevant vocabulary under a unified topic. However, an analysis of the highest-scoring words from Topics 4 and 6 reveals that words of different grammatical forms often convey essentially the same meaning. Both ‘recycling’ and ‘recycle’ in the text reflect the concept of resource recovery, while ‘sustainability’ and ‘sustainable’ repeatedly express the notion of environmental sustainability. We believe that by leveraging the strengths of large language models, including part-of-speech normalisation and selective word comparison, BERTopic can offer researchers a more comprehensive vocabulary pertaining to each topic.
The frequencies and corresponding times of the seven themes are presented in Figure 4. Topic 1 shows a growing trajectory, rising significantly starting around 2017. This upgrade is timed to coincide with key international developments such as the 2016 Paris Agreement, highlighting the growing focus in the UK news on the synergies between renewable energy adoption and carbon emissions. Topic 15 experienced a surge in frequency around 2019. The topic continues to gain visibility, and this rise may be related to increased public concern about chemical pollution incidents and environmental health. The increase in frequencies of all topics post-2015 suggests an overall growing public and journalistic interest in environmental issues. This period seems to be pivotal in shaping the public sentiment on sustainability and environmental topics.
Figure 5 presents a heatmap of the similarity matrix for topics generated by the BERTopic model on The Guardian dataset. These seven topics correspond to the ones shown in Figure 3. Deeper shades represent higher cosine similarity, whereas lighter shades indicate lower similarity. Since we get more topics in the Guardian dataset than in the Reddit and Twitter dataset, we believe the heatmap in Figure 5 facilitates a comparative analysis of the interrelationships among the topics.
According to this scatter plot, we can notice that the cluster of Topic 14 (pink colour) is somewhat separated from the rest of the topic clusters. This means that the news around electric vehicles in The Guardian is more of a dedicated discussion to this area, focusing on key industry players such as Tesla. While this topic may not be at the centre of most discussions in The Guardian news, it can be a good representation of the impact of innovation trends, CE being carried out in transport. Additionally, although the Topic 15 (grey colour) is similar to other clusters, it has a smaller size which may reflect the specificity of discussions around toxic chemicals. The size of clusters in this topic shows there is a specialised and smaller number of news about chemical term and EPA in The Guardian dataset.
In the field of topic modelling, word clouds serve as a visual heuristic method, displaying terms that occupy central nodes in the discourse network [59]. Figure 7 shows the topic vocabulary graph cloud generated by BERTopic on The Guardian dataset. The visual prominence of terms such as ‘sustainability’, ‘sustainable’, ‘recycling’, and ‘energy’ within the cloud is indicative of their pivotal role in the corpus, highlighting the dataset’s focus on environmental conservation, renewable energy, and the principles of the CE. These terms underscoring the centrality of sustainable energy in CE are captured within The Guardian news. Surrounding these core terms are secondary, yet substantial, concepts such as ‘ev’, ‘electric’, and ‘vehicle’, which, given their relative visual weight (size), suggest substantive topics that concern technological advancements in transportation and the transition. Moreover, the adjacency of concepts such as ‘ethical’, ‘consumer’, ‘growth’, and ‘future’ alongside ’sustainability’ within the visualisation elucidates an interplay of notions.
4.3 Professional Sources: Reddit
The word cloud (Figure 11) not only captures the direction of Reddit user’s discussions about CE but also provides insight into the collective focal points of the community, thereby reflecting the broader public engagement and sentiment towards sustainability.
4.4 Public Sources: Twitter
@@
The line graph (Figure 13) displays the fluctuation frequencies of five distinct topics generated by the CorEx model based on 2023 Twitter dataset. These topics exhibit significant frequency peaks over time, which may correspond to specific events that triggered Twitter discussions or heightened public sentiment. Topic 2 shows less volatility compared to others but experiences a sharp increase in discussion volume. This likely relates to an event connected with ‘recycle’, temporarily bringing this topic to the forefront on Twitter. Notably, Topic 4 demonstrates a very pronounced peak, exceeding the activity of other topics during the same period. This outlier could represent a major event or breakthrough in sustainability towards the end of March 2023, or an activity that briefly dominated discussions. Figure 13 reflects Twitter’s reactive nature as a platform for public discourse. A temporal analysis of topic frequencies can provide deeper insights into how certain themes gain traction within the social media community, highlighting the interplay between global events and public sentiment.
@@ 5.1 Conclusion
In the context of sustainable transformation, this study aims to uncover public sentiment of the CE by analysing data from various sources, including social media platforms Twitter and Reddit, as well as the Guardian. Utilising topic modelling, this paper explores the multi-perspective and changing perceptions of the CE among the public, professionals, and officials. Three topic models(LDA, CorEx, and BERTopic) were employed. The findings of this study reveal the public’s growing concern for environmental issues, highlighting widespread focus on topics such as sustainability, recycling, and CE. Furthermore, the dynamic evolution of these topics over time indicates that public interest in the CE is not only increasing but also diversifying. This reflects a deepening engagement with the CE, influenced by global and regional environmental events. Through systematic experiments and optimal hyperparameter tuning of the models, we have successfully captured shifts in public sentiment and opinions, thereby providing a comprehensive view of public concerns regarding the CE. Overall, this study contributes to both academic and practical understandings of public perceptions in the context of the CE. By mapping the cognitive pathways and concerns of different segments of society, this paper aids policymakers and businesses in adapting strategies and communications to promote a more sustainable future.
@@ 5.2 Policy Implications
Our study also offers important policy implications. Firstly, this study highlights the growing public sentiment and concern regarding sustainability, recycling, and the CE. To leverage this trend, policymakers should implement targeted education programs to deepen public knowledge and understanding of the CE [60]. These programs can emphasise the benefits of circular practices, illustrating how individuals can contribute to environmental protection through actions such as recycling and choosing sustainable products [61]. Additionally, community engagement programs can be designed to involve local residents in CE initiatives, thereby fostering public awareness and responsibility for environmental stewardship. It is essential that these policies are tailored for different regions and demographic groups to maximise their reach and impact.
Furthermore, the diverse public sentiment on various aspects of the CE, such as emissions reduction and recycling, demonstrates strong support for businesses that adopt CE practices [45]. Policymakers should consider creating incentives for businesses that integrate CE principles into their operations. For example, companies that reduce waste, use recycled materials, or develop innovative products to facilitate a transition to the CE could receive tax exemptions, subsidies, or other incentives [62]. Additionally, policies can be adjusted to encourage companies to take greater responsibility in advancing the CE, ensuring that businesses operate within a clear regulatory framework that promotes sustainable development.
This study also revealed significant and growing public sentiment regarding recycling and waste management. In response, policymakers should therefore enhance existing regulations to establish a more stringent waste management policies and improved recycling processes [63]. Measures could include setting higher recycling targets for government authorities, imposing stricter controls on industrial waste, and providing incentives for companies that significantly reduce waste production. Expanding waste management infrastructure can also help meet the growing public demand for efficient recycling services [64]. Moreover, policies should encourage the development of new technologies that enhance recycling efficiency and waste reprocessing capabilities, ultimately making the CE more effective and accessible.
@@ 5.3 Future Research Avenue
We will refine the granularity of our topic models by utilising more advanced hyperparameter tuning algorithms to optimise models with higher coherence NPMI scores. Furthermore, we will conduct deeper explorations based on the BERTopic model, which has shown optimal performance in both datasets used in this study. This includes investigating multimodal approaches that incorporate textual data as well as visual and auditory content, offering a more comprehensive analysis of public sentiment. This might involve integrating Transformer-based architectures [65] with computer vision techniques to analyse images and videos related to the CE on social media, thereby providing a richer and more thorough understanding of public discourse. This integrated approach not only enhances the precision of sentiment and thematic detection but also captures a broader array of data points for sentiment analysis, including emotional expressions and visual symbols. Longitudinal and cross-platform studies are crucial for observing changes in the emotions of different user groups over time, thereby gaining deeper insights into the dynamic shifts in public engagement with sustainability issues.
@@ Acknowledgements
This research is funded by Digital Circular Electrochemical Economy (DCEE) EP/V042432/1, and the UK Research and Innovation (UKRI) Interdisciplinary Centre for Circular Chemical Economy EP/V011863/1.

### v2

@@ 4.5.1 SOO
In conclusion, BERTopic provided a deep extraction of CE topics through a language model (BERT), CorEx provided a unique separation of topics (highest coherence score), and LDA captured the underlying concepts of CE on Twitter, all of which provide valuable insight into our understanding of the public attention on the CE. In the next three subsections, we will visualise and discuss the topics generated by the best models on each dataset in more depth.
@@
Visualisation and Analysis
@@
The line graph (Figure 15) displays the fluctuation frequencies of five distinct topics generated by the CorEx model based on 2023 Twitter dataset. These topics exhibit significant frequency peaks over time, which may correspond to specific events that triggered Twitter discussions or heightened public attention. Topic 2 shows less volatility compared to others but experiences a sharp increase in discussion volume. This likely relates to an event connected with ‘recycle’, temporarily bringing this topic to the forefront on Twitter. Notably, Topic 4 demonstrates a very pronounced peak, exceeding the activity of other topics during the same period. This outlier could represent a major event or breakthrough in sustainability towards the end of March 2023, or an activity that briefly dominated discussions. Figure 15 reflects Twitter’s reactive nature as a platform for public discourse. A temporal analysis of topic frequencies can provide deeper insights into how certain topics gain traction within the social media community, highlighting the interplay between global events and public attention.
@@
4.5.2 MOO
The topics generated by BERTopic on the Twitter dataset, optimised using MOO as shown in Table 17. For instance, Topic 0 includes terms like ‘circulareconomy’, ‘recycling’ and ‘circular’, directly addressing key aspects of resource efficiency and waste management central to CE. The frequency of these terms indicates that users on Twitter are actively engaging with discussions about recycling and resource reuse, reflecting the platform’s role in disseminating information and promoting environmental advocacy.
Topic 2 delves into the specifics of sustainable packaging and recycling, with terms such as ‘sustainablepackage’ and ‘recycledmaterial’ demonstrating how MOO can highlight niche topics within broader CE discussions. This kind of granular analysis is particularly useful for identifying targeted conversations about sustainability in industry, especially in relation to packaging materials.
Table 17 shows the diversity of topics, ranging from charity-related support initiatives (Topic 3) to political and legislative discussions on plastic waste (Topic 4). The ability of MOO to uncover such a wide array of topics underscores its value in capturing the breadth of public discourse on CE. Overall, MOO allows for a more detailed extraction of the evolving narratives on platforms like Twitter, revealing both grassroots actions and industry-specific concerns.
On the Twitter platform, the topic model generated by SOO optimisation exhibits relatively good consistency, especially with a smaller number of topics (e.g., 10), which focus on key subjects such as ‘emission’, ‘carbon’, and ‘recycling’. However, because SOO only optimises coherence, the generated topics lack diversity and fail to fully reflect the range of user discussions on Twitter, as well as the platform’s characteristic short text format.
MOO increases the number of topics (8) and improves the diversity score, capturing more discussions related to the circular economy. Although MOO’s consistency is slightly lower than that of SOO, the significantly reduced perplexity suggests that MOO better handles content on short-text platforms like Twitter. Additionally, MOO generates topics that cover a wider variety of subjects, such as electric vehicles (EVs) and environmental policies, making it more adaptable to the content diversity and brevity typical of Twitter. Therefore, MOO is better suited for platforms like Twitter, where topics are diverse and discussions are more dispersed.
Table 17: The topic terms obtained from the BERTopic model with the best hyperparameter values on the Twitter dataset (MOO).
Topic Topic Term
0 circulareconomy, recycling, recycle, circular, sustainable, wastemanagement, sustainability, circularity, join, 2023
1 circulareconomy, circularity, circular, sustainabledevelopment, economic, future, economy, sustainable, consumption, reuse
2 sustainablepackage, recycledmaterial, recyclability, sustainable, sustainability, recycling, recyclable, plasticpollution, flexiblepackage, recyclingbusiness
3 charitynee, support, charityneed, supportlocal, recycle, charity, need, goodwill, goodwillfind, donate
4 plasticwaste, wastetrade, endwastecolonialism, wasteshipment, waste, plasticpollution, recycling, eu, circulareconomy, ban
5 c3650, cisco, gigabit, c6524gs, dl360, ethernet, port, 3650, server, uplink
6 cisco1921, cisco, router, processor, performace, forsale, ethernet, 12000, c887vam, circulareconomy
7 x450e, x450, x450a, x460, extreme, summit48si, poe, 24p, forsale, 24pt
@@ 5.1 Conclusion
In the context of sustainable transformation, this research aims to uncover public attention to the CE by analysing data from various sources, including social media platforms Twitter and Reddit, as well as the Guardian. By utilising topic modelling, this paper examines the multi-perspective and changing perceptions of the CE among the public, professionals, and officials. Three topic models (LDA, CorEx, and BERTopic) were employed. Our findings reveal a growing public concern for environmental issues, with a strong focus on sustainability, recycling, and circular practices. Furthermore, the dynamic evolution of these topics over time indicates that public interest in the CE is not only increasing but also diversifying. This reflects a deepening engagement with the CE, influenced by global and regional environmental events. Through systematic experiments and twin hyperparameter optimisation of the models, we effectively captured shifts in public attention and opinions, thereby providing a comprehensive view of societal concerns regarding the CE. Overall, this study contributes to both academic and practical understandings of public perceptions of the CE, mapping the cognitive pathways and concerns across different societal segments, thus offering valuable insights for policymakers and businesses to adapt strategies and communications for a more sustainable future.
This research compares and analyses the results of SOO (single-objective optimisation) and MOO (multi-objective optimisation) on CE-related topics across different platforms (The Guardian, Reddit, Twitter), highlighting the differences between these optimisation methods in topic model generation. SOO generates topics with high semantic relevance by maximising topic coherence, but it lacks diversity and adaptability, especially on platforms with dispersed discussions (such as Reddit and Twitter), where the number of generated topics is small and the information coverage is limited. In contrast, MOO produces a more diverse set of topics by optimising coherence, diversity, and perplexity simultaneously, demonstrating stronger adaptability on platforms with more complex and short-form texts. MOO not only enhances the model’s ability to capture diverse discussion content but also significantly reduces perplexity while maintaining topic coherence, enabling it to handle unseen texts more robustly. In summary, MOO’s application across different platforms proves it to be more suitable for analysing CE topics, providing a more comprehensive understanding of public concerns and discussions.
@@ 5.2 Policy Implications
Our study also offers important policy implications. First, the growing public attention and concern regarding sustainability, recycling, and the CE underscores the need for targeted education programs to enhance public knowledge of CE principles [69]. Such programs should emphasise the benefits of circular practices, illustrating how individuals actions, like recycling and choosing sustainable products, contribute to environmental protection [70]. Additionally, community engagement can be designed to involve local residents in CE initiatives, thereby fostering public awareness and responsibility for environmental stewardship. Tailoring these policies to different regions and demographic groups will help maximise their reach and impact.
Furthermore, the study highlights strong public support for businesses that adopt CE practices (e.g., emissions reduction, reduce waste, use recycled materials, developing innovative circular products) [53]. Policymakers should consider creating incentives (e.g., tax exemptions, subsidies) for businesses that integrate CE principles into their operations [71]. Additionally, policies should encourage companies to take greater responsibility in advancing the CE, ensuring that businesses operate within a clear regulatory framework that promotes sustainable development.
Given the significant and growing public attention to recycling and waste management, policymakers should enhance existing regulations to establish more stringent waste management policies and improved recycling processes [72]. Measures could include setting higher recycling targets for government authorities, imposing stricter controls on industrial waste, and providing incentives for companies that significantly reduce waste production. Expanding waste management infrastructure will also be crucial to meeting growing public demand for efficient recycling services [73]. Moreover, policies should support the development of new technologies that improve recycling efficiency and waste reprocessing capabilities, ultimately making CE practices more accessible and effective.
@@ 5.3 Future Research Avenue
Future research will focus on conducting deeper explorations based on the BERTopic model, which has shown optimal performance in both datasets used in this study. This includes investigating multimodal approaches that incorporate textual data as well as visual and auditory content, offering a more comprehensive analysis of public attention. This might involve integrating Transformer-based architectures [74] with computer vision techniques to analyse images and videos related to the CE on social media, thereby providing a richer and more thorough understanding of public discourse. This integrated approach not only enhances the precision of attention and thematic detection but also captures a broader array of data points for attention analysis, including emotional expressions and visual symbols. Longitudinal and cross-platform studies are crucial for observing changes in the emotions of different user groups over time, thereby gaining deeper insights into the dynamic shifts in public engagement with sustainability issues.
@@ Acknowledgements
This research is funded by Digital Circular Electrochemical Economy (DCEE) EP/V042432/1, and the UK Research and Innovation (UKRI) Interdisciplinary Centre for Circular Chemical Economy EP/V011863/1 and EP/V011863/2.
Appendix A
Table 18: 53 keywords related to circular economy for data scraping. Based on the existing literature, this study first summarises the keywords related to the circular economy [1, 6, 60]. Then, these keywords are expanded using the word embedding model to develop a comprehensive circular economy dictionary. As a novel variable quantification method, the word embedding model improves the accuracy of variable measurement and enhances the robustness of empirical research results [75]. This model employs neural networks to perform deep analysis on a large volume of economic and financial texts, generating a word similarity model that can identify and train similar words. The similarity dictionary produced by this model allows variables to be measured more comprehensively and objectively [75].
sustainable development energy saving and environmental protection energy saving
green economy eco-design energy conservation
clean manufacturing reuse zero waste
renewable energy reducing carbon footprint renewable resources
waste reduction low carbon closed-loop system
energy efficiency green manufacturing environment friendly
green procurement recycling environmentally friendly
green natural capital biomimicry
social responsibility environmental economics efficient use
regeneration cycle sustainable supply chain industrial ecology
durability environmental impact assessment resource efficiency
sharing economy eco-friendly products save resources
green low carbon clean production sustainable consumption
eco-efficiency resource saving clean
environmental protection clean energy biodegradable material
extended producer responsibility ecological civilization environmental governance
zero emission resource recovery remanufacturing
waste management green ecology